### v1

# 37N35, 65F55, 90C22, 90C59, 93B52

Abstract
Learning stabilizing controllers from data is an important task in engineering applications; however, collecting informative data is challenging because unstable systems often lead to rapidly growing or erratic trajectories. In this work, we propose an adaptive sampling scheme that generates data while simultaneously stabilizing the system to avoid instabilities during the data collection. Under mild assumptions, the approach provably generates data sets that are informative for stabilization and have minimal size. The numerical experiments demonstrate that controller inference with the novel adaptive sampling approach learns controllers with up to one order of magnitude fewer data samples than unguided data generation. The results show that the proposed approach opens the door to stabilizing systems in edge cases and limit states where instabilities often occur and data collection is inherently difficult.
keywords:
nonlinear systems, feedback stabilization, context-aware learning, informative data, data-driven control
[…]
We introduce two multi-fidelity variants an adaptive sampling scheme for the data-driven construction of stabilizing feedback controllers for nonlinear dynamical systems. The proposed method generates informative data while simultaneously stabilizing the system to avoid instabilities during the data collection, which otherwise could render the collected data meaningless.
1 Introduction
The stabilization of dynamical systems is an essential building block in many applications ranging from the control of power networks to chemical reactors to fluid structure interactions. When explicit model descriptions are unavailable, data has to be used for the design of suitable stabilizing controllers. However, the collection of data from unstable systems is a challenging task since the dynamical instabilities quickly render the collected data uninformative. In this work, we consider the design of informative data sets for the task of system stabilization via an adaptive data sampling scheme.
Many approaches have been considered for the task of data-driven control ranging from model-free parameter tuning approaches [6, 9, 19, 28] to scientific machine learning and reinforcement learning methods [14, 27, 17, 44]. With the rise of data-driven reduced-order modeling, a two-step procedure has been established in which first, a model of the underlying dynamics is learned, before this model is used for the design of controllers via classical model-based approaches. Popular data-driven modeling methods for the task of control include among others the dynamic mode decomposition and operator inference [30, 36, 26, 24, 34, 3], sparse identification methods [5, 29], and the Loewner framework [21, 31, 25, 32, 1, 11]. In recent years, the concept of data informativity has emerged as an effective tool for solving a variety of data-driven tasks including system stabilization [8, 39, 43, 42]. This framework is well suited for the development of new data-driven approaches as it is based on the question of how much data is needed to fulfill a considered task. While there is such a variety of different approaches for designing controllers from given data, most of these methods assume that the data contain enough information about the system to solve the task at hand. Here, we aim to design such informative data sets.
A significant component in the data sampling for dynamical processes are the external inputs of the systems. Classically, these are chosen to be random signals. Such inputs satisfy with probability one the persistent excitation condition that guarantees, for example, upper bounds on the sample complexity for system identification [45]. While this choice of inputs allows to eventually identify all information about the system, it often is too generic to achieve small sampling numbers for the particular task of designing stabilizing controllers for a particular given system. In fact, it has been shown in [43] that there exist significantly smaller data sets that are informative for the task of system stabilization. In this work, we propose to construct such informative data sets by iteratively updating the input signals for the system based on a sequence of low-dimensional stabilizing controllers over subspaces computed via low-dimensional controller inference. These subspaces are expanded using the sampled data until the full unstable dynamics of the underlying system are stabilized. The main motivation for this approach is that the sample complexity of low-dimensional controller inference only scales with the dimension of the subspace such that only few data samples are needed to stabilize the system in the low-dimensional subspace, as we will show. Besides a lower sample complexity for constructing stabilizing controllers, the proposed adaptive data generation suppresses the dynamical instabilities during the sampling process, ensuring informative data for stabilization.
The remainder of this manuscript is organized as follows. In Section 2, we remind the reader of basic concepts needed for the stabilization of dynamical systems and the data informativity framework for learning stabilizing controllers. We derive the adaptive data sampling scheme in Section 3 based on theoretical results that outline suitable input signals, and we discuss convergence of the method as well as extensions for stable numerical computations. Numerical experiments are performed in Section 4 including the stabilization of a power network and the laminar flow behind an obstacle. The paper is concluded in Section 5.
2 Preliminaries
2.1 Stabilization of dynamical systems
Let $(x^{\operatorname{n}}(t))_{t\geq 0}$ be a process that can be controlled with inputs $(u^{\operatorname{n}}(t))_{t\geq 0}$, where $x^{\operatorname{n}}(t)\in\mathbb{R}^{N}$ and $u^{\operatorname{n}}(t)\in\mathbb{R}^{p}$ are state and control at time $t$, respectively. For feasible initial conditions $x^{\operatorname{n}}(0)\in\mathcal{X}_{0}^{\operatorname{n}}\subseteq\mathbb{R}^{N}$, the dynamics of $x^{\operatorname{n}}(t)$ are described in continuous time by systems of ordinary differential equations
$\frac{\operatorname{d}}{\operatorname{d}t}x^{\operatorname{n}}(t)=f(x^{\operatorname{n}}(t),u^{\operatorname{n}}(t)),\quad t\geq 0,$ (1)
with the time evolution function $f\colon\mathbb{R}^{N}\times\mathbb{R}^{p}\to\mathbb{R}^{N}$, which can be nonlinear in both arguments. In discrete time, processes are governed by systems of difference equations
$x^{\operatorname{n}}(t+1)=f(x^{\operatorname{n}}(t),u^{\operatorname{n}}(t)),\quad t\in\mathbb{N}\cup\{0\}\,.$ (2)
We are interested in stabilizing processes towards their steady states. A steady state $(\bar{x},\bar{u})$ is an equilibrium of the time evolution function such that in the continuous-time case eq. 1 it holds
$0=f(\bar{x},\bar{u}),$ (3)
and for discrete-time systems eq. 2 it holds
$\bar{x}=f(\bar{x},\bar{u}).$ (4)
Steady states are called unstable if trajectories do not converge to $(\bar{x},\bar{u})$ for initial conditions $x^{\operatorname{n}}(0)$ in an arbitrarily small neighborhood about $(\bar{x},\bar{u})$. The task we consider is to construct a state-feedback controller $K\in\mathbb{R}^{p\times N}$ such that there exists an $\epsilon>0$ so that for all initial conditions $x^{\operatorname{n}}(0)\in\mathcal{X}_{0}^{\operatorname{n}}$ with $\lVert x^{\operatorname{n}}(0)-\bar{x}\rVert\leq\epsilon$, the control signal given by
$u^{\operatorname{n}}(t)=K(x^{\operatorname{n}}(t)-\bar{x})+\bar{u}$ (5)
leads to $x^{\operatorname{n}}(t)\to\bar{x}$ in the limit $t\to\infty$. Such a controller $K$ is then denoted as locally stabilizing.
2.2 Controlling systems from data
In this work, a model of the system dynamics in terms of the right-hand side function $f$ in eq. 1 or eq. 2 is unavailable in the sense that $f$ is not given in closed form. Instead, we assume that $f$ can be evaluated for suitable states and control inputs. For either system type eq. 1 or eq. 2, this leads to data triplets $(U^{\operatorname{n}},X^{\operatorname{n}},Y^{\operatorname{n}})$ of the form
$U^{\operatorname{n}}$ $=\begin{bmatrix}u^{\operatorname{n}}(t_{1})&u^{\operatorname{n}}(t_{2})&\ldots&u^{\operatorname{n}}(t_{T})\end{bmatrix},$ (6a)
$X^{\operatorname{n}}$ $=\begin{bmatrix}x^{\operatorname{n}}(t_{1})&x^{\operatorname{n}}(t_{2})&\ldots&x^{\operatorname{n}}(t_{T})\end{bmatrix},$ (6b)
$Y^{\operatorname{n}}$ $=\begin{bmatrix}f(x^{\operatorname{n}}(t_{1}),u^{\operatorname{n}}(t_{1}))&f(x^{\operatorname{n}}(t_{2}),u^{\operatorname{n}}(t_{2}))&\ldots&f(x^{\operatorname{n}}(t_{T}),u^{\operatorname{n}}(t_{T}))\end{bmatrix},$ (6c)
at some discrete time points $t_{1},t_{2},\ldots,t_{T}$. These time points do not need to be ordered or have a particular distance to each other and may even be identical.
For continuous-time systems, the columns of $Y^{\operatorname{n}}$ in eq. 6c are the derivatives of eq. 6b with respect to time such that
$Y^{\operatorname{n}}=\begin{bmatrix}\frac{\operatorname{d}}{\operatorname{d}t}x^{\operatorname{n}}(t_{1})&\frac{\operatorname{d}}{\operatorname{d}t}x^{\operatorname{n}}(t_{2})&\ldots&\frac{\operatorname{d}}{\operatorname{d}t}x^{\operatorname{n}}(t_{T})\end{bmatrix}.$
The simulation of trajectories of continuous-time systems eq. 1 typically yields a discretization error. However, the considered data setup is independent of this type of error since only evaluation pairs of the time evolution function are used and no exact, continuous trajectories over multiple time steps are needed. Note that if the time derivatives are only approximated using the states of consecutive time steps, the time discretization as well as approximation errors play a role. However, these errors can typically be controlled by the choice of the time-step size and discretization schemes. We do not discuss this particular case in this work.
In the case of discrete-time systems eq. 2, the evaluation of the evolution function $f$ amounts to single forward time steps on trajectories such that $Y^{\operatorname{n}}$ in eq. 6c simplifies to
$Y^{\operatorname{n}}=\begin{bmatrix}x^{\operatorname{n}}(t_{1}+1)&x^{\operatorname{n}}(t_{2}+1)&\ldots&x^{\operatorname{n}}(t_{T}+1)\end{bmatrix}.$
A classical choice for eq. 6 in this case is single trajectory data resulting in $Y^{\operatorname{n}}$ being $X^{\operatorname{n}}$ shifted by a single time step such that
$U^{\operatorname{n}}$ $=\begin{bmatrix}u^{\operatorname{n}}(0)&u^{\operatorname{n}}(1)&\ldots&u^{\operatorname{n}}(T-1)\end{bmatrix},$
$X^{\operatorname{n}}$ $=\begin{bmatrix}x^{\operatorname{n}}(0)&x^{\operatorname{n}}(1)&\ldots&x^{\operatorname{n}}(T-1)\end{bmatrix},$
$Y^{\operatorname{n}}$ $=\begin{bmatrix}x^{\operatorname{n}}(1)&x^{\operatorname{n}}(2)&\ldots&x^{\operatorname{n}}(T)\end{bmatrix};$
see, for example, [43, 42, 39, 8, 30, 5].
2.3 Data informativity and controller inference
In contrast to traditional approaches, which train some chosen parametrization of $f$ using eq. 6 to infer a model describing the system dynamics, we are going to directly learn a state feedback controller $K$ that stabilizes the unknown system described by $f$. This has been shown to allow for significant reductions in terms of data needed to solve the task [43, 42, 39, 8]. A general concept for data-driven tasks that has been developed in recent years is data informativity [45, 38]. Underlying is the question, how much data of the form eq. 6 is needed to achieve a task. This circumvents the need of identifying the true underlying system dynamics and allows to formulate sample complexity results in terms of the task at hand.
Necessary and sufficient conditions for data informativity for the stabilization of linear dynamical systems by state-feedback have been developed in [39, 8] and extended to low-dimensional subspaces in [43]. A relation to nonlinear systems has been outlined in [42], which we will quickly recap in the following. For analytic models $f$ in eqs. 1 and 2, the local dynamics about the steady state $(\bar{x},\bar{u})$ can be described by a linearization of the form
$f(x^{\operatorname{n}}(t),u^{\operatorname{n}}(t))$ $=f(\bar{x},\bar{u})+J_{\operatorname{x}}(x^{\operatorname{n}}(t)-\bar{x})+J_{\operatorname{u}}(u^{\operatorname{n}}(t)-\bar{u})$ (7)
$\quad{}+{}\mathcal{O}(\lVert x^{\operatorname{n}}(t)-\bar{x}\rVert^{2})+\mathcal{O}(\lVert u^{\operatorname{n}}(t)-\bar{u}\rVert^{2}),$
where the Jacobians of $f$ with respect to the states $x$ and inputs $u$ evaluated in the steady state $(\bar{x},\bar{u})$ are given by
$J_{\operatorname{x}}=\nabla_{\operatorname{x}}f(\bar{x},\bar{u})\quad\text{and}\quad J_{\operatorname{u}}=\nabla_{\operatorname{u}}f(\bar{x},\bar{u}).$
The deviation from the desired steady state $x(t)=x^{\operatorname{n}}(t)-\bar{x}$ is then approximated by the linear systems
$\frac{\operatorname{d}}{\operatorname{d}t}x(t)=J_{\operatorname{x}}x(t)+J_{\operatorname{u}}u(t)\quad\text{and}\quad x(t+1)=J_{\operatorname{x}}x(t)+J_{\operatorname{u}}u(t),$ (8)
for continues and discrete-time systems, respectively, with $u(t)=u^{\operatorname{n}}(t)-\bar{u}$ and the initial condition $x(0)\in\mathcal{X}^{\operatorname{n}}_{0}-\bar{x}$. The data triplets $(U,X,Y)$ corresponding to eq. 8 are related to eq. 6 as approximations to columnwise shifts of the data given by
$U$ $=U^{\operatorname{n}}-\bar{u}=\begin{bmatrix}u^{\operatorname{n}}(t_{1})-\bar{u}&\ldots&u^{\operatorname{n}}(t_{T})-\bar{u}\end{bmatrix},$ (9a)
$\widetilde{X}$ $=X^{\operatorname{n}}-\bar{x}=\begin{bmatrix}x^{\operatorname{n}}(t_{1})-\bar{x}&\ldots&x^{\operatorname{n}}(t_{T})-\bar{x}\end{bmatrix},$ (9b)
$\widetilde{Y}$ $=Y^{\operatorname{n}}-f(\bar{x},\bar{u})$ (9c)
$=\begin{bmatrix}f(x^{\operatorname{n}}(t_{1}),u^{\operatorname{n}}(t_{1}))-f(\bar{x},\bar{u})&\ldots&f(x^{\operatorname{n}}(t_{T}),u^{\operatorname{n}}(t_{T}))-f(\bar{x},\bar{u})\end{bmatrix}$
such that
$X=\widetilde{X}+\mathcal{O}(\lVert\widetilde{X}\rVert^{2})\quad\text{and}\quad Y=\widetilde{Y}+\mathcal{O}(\lVert\widetilde{Y}\rVert^{2})+\mathcal{O}(\lVert U\rVert^{2}).$ (10)
The linear controller $K$ in eq. 5 can be designed from idealized data triplets $(U,X,Y)$ of linear systems eq. 8 via data informativity conditions as outlined in the following proposition.
Proposition 1 (Data informativity for stabilization [39, 43]).
Let $(U,X,Y)$ be a data triplet sampled from a linear state-space model. The data triplet is informative for stabilization by feedback if and only if one of the following equivalent statements hold:
  • (a)
    The matrix $X$ has full row rank and there exists a right inverse $X^{\dagger}$ of $X$ such that $YX^{\dagger}$ has only stable eigenvalues. A stabilizing controller is then given by $K=UX^{\dagger}$.
  • (b)
    There exists a matrix $\Theta\in\mathbb{R}^{T\times N}$ such that if the sampled model is continuous in time
    $X\Theta>0\quad\text{and}\quad Y\Theta+\Theta^{\mkern-1.5mu\mathsf{T}}Y^{\mkern-1.5mu\mathsf{T}}<0$ (11)
    hold, or if the sampled model is discrete in time
    $X\Theta=(X\Theta)^{\mkern-1.5mu\mathsf{T}}\quad\text{and}\quad\begin{bmatrix}X\Theta&Y\Theta\\ (Y\Theta)^{\mkern-1.5mu\mathsf{T}}&X\Theta\end{bmatrix}>0$ (12)
    hold. A stabilizing controller is then given by $K=U\Theta(X\Theta)^{-1}$.
The eigenvalue criterion in part (a) of Proposition 1 must be treated differently depending on the type of model that is considered similar to part (b). A matrix has only stable eigenvalues if in the continuous-time case the real part of all eigenvalues is negative, or if in the discrete-time case the absolute value of all eigenvalues is smaller than one.
Following Proposition 1, we call data informative for stabilization if there exists a controller $K$ that stabilizes all linear state-space models given by matrices $A\in\mathbb{R}^{N\times N}$ and $B\in\mathbb{R}^{N\times p}$ that describe the data $(U,X,Y)$ via $Y=AX+BU$. In other words, the data contain enough information to solve the stabilization problem.
The necessary condition for data informativity, namely that $X$ has full row rank, can only be satisfied if the subspace of reachable system states is the full state space $\mathbb{R}^{N}$. In the case that system states span only a subspace $\mathcal{V}$ of dimension $r<N$, the classical concept of data informativity (Proposition 1) cannot be applied anymore. A low-dimensional data informativity concept for the task of stabilization has been developed in [43]. Thereby, the conditions in Proposition 1 are not considered in the full $N$-dimensional state space but only in the $r$-dimensional subspace $\mathcal{V}$ using a projection.
Proposition 2 (Low-dimensional data informativity [43]).
Let $(U,X,Y)$ be a data triplet sampled from a linear state-space model of dimension $N$ and let $\mathcal{V}$ be the $r$-dimensional subspace spanned by all trajectories of the model such that for an orthogonal basis matrix $V$ of $\mathcal{V}$, the state satisfies $x(t)=V\hat{x}(t)$ for all $t\geq 0$, where $\hat{x}(t)\in\mathbb{R}^{r}$ is the state of a linear $r$-dimensional model. The data $(U,X,Y)$ are informative for stabilization over $\mathcal{V}$ if and only if $(U,\widehat{X},\widehat{Y})$ satisfies Proposition 1, with
$\widehat{X}=V^{\mkern-1.5mu\mathsf{T}}X\quad\text{and}\quad\widehat{Y}=V^{\mkern-1.5mu\mathsf{T}}Y.$
A stabilizing controller is then given by $K=\widehat{K}V^{\mkern-1.5mu\mathsf{T}}$, where $\widehat{K}$ is a controller that stabilizes all linear models that describe the projected data $(U,\widehat{X},\widehat{Y})$.
Proposition 2 allows to design a stabilizing controller $K$ for an $N$-dimensional model using only $r$ many data samples. The state equality in Proposition 2 is necessary to give a guarantee for the stabilization property; however, in practice it has been observed that already an approximate relation $x(t)\approx V\hat{x}(t)$ for some dimension $r<n_{\min}$, where $n_{\min}$ is the intrinsic system dimension, is sufficient to design a stabilizing controller via Proposition 2. The application of the stabilization theory to nonlinear systems is possible via system linearization based on eq. 7. Under the assumption that the model $f$ is analytic, it holds that for $x^{\operatorname{n}}(t)$ staying close enough to the steady state $\bar{x}$, a controller that stabilizes the corresponding linearization eq. 8 also stabilizes the nonlinear system; see [42, Prop. 3.1]. With a similar argument, we can use the shifted data $(U,\widetilde{X},\widetilde{Y})$ as approximately linear data because the relation eq. 10 holds and design a stabilizing controller from Propositions 1 and 2. The resulting method for low-dimensional controller inference of nonlinear systems is summarized in Algorithm 1. The algorithm operates on the data from the nonlinear model eq. 1 or eq. 2. We call the controller inferred by Algorithm 1 stabilizing over the subspace $\mathcal{V}$ with basis $V$ as it is stabilizing all nonlinear systems with linearizations that describe the shifted and projected data
$U=U^{\operatorname{n}}-\bar{u},\quad\widehat{X}=V^{\mkern-1.5mu\mathsf{T}}(X^{\operatorname{n}}-\bar{x})\quad\text{and}\quad\widehat{Y}=V^{\mkern-1.5mu\mathsf{T}}(Y^{\operatorname{n}}-f(\bar{x},\bar{u})).$
Note that the subspace $\mathcal{V}$ can be chosen arbitrarily. However, a typically suitable choice, which satisfies $x(t)\approx V\hat{x}(t)$, can be computed directly from the given data $X^{\operatorname{n}}-\bar{x}$ using, for example, the singular value decomposition.
1
Input: High-dimensional data triplet $(U^{\operatorname{n}},X^{\operatorname{n}},Y^{\operatorname{n}})$, orthogonal basis matrix $V$.
2
Output: State-feedback controller $K$.
3
4Compute the reduced and shifted data triplet $(U,\widehat{X},\widehat{Y})$ via
$U=U^{\operatorname{n}}-\bar{u},\quad\widehat{X}=V^{\mkern-1.5mu\mathsf{T}}(X^{\operatorname{n}}-\bar{x})\quad\text{and}\quad\widehat{Y}=V^{\mkern-1.5mu\mathsf{T}}(Y^{\operatorname{n}}-f(\bar{x},\bar{u})).$
5
6Infer a low-dimensional stabilizing feedback $\widehat{K}=U\Theta(\widehat{X}\Theta)^{-1}$ for $(U,\widehat{X},\widehat{Y})$ solving either eq. 11 or eq. 12 for the unknown $\Theta$.
7
8Lift the inferred controller $\widehat{K}$ to the high-dimensional space via $K=\widehat{K}V^{\mkern-1.5mu\mathsf{T}}$.
Algorithm 1 Low-dimensional controller inference.
2.4 Problem statement
A crucial point in the design of controllers from data is the quality of the data. If we have the possibility to influence the generation of data by exciting a system using initial values and control signals, the question is how to generate an informative data set for stabilization in the sense of eq. 6. Classical choices from system identification [4, 37, 45] such as randomized input signals to generate state trajectories are generic and do not take the specific task of stabilization into account. Furthermore, using random input signals provides no way to steer the collection process and thus can lead to unbounded growth, rendering computations with the generated data numerically infeasible, and limit cycle behavior, in which case system states are repeated and therefore the information content of the data is limited.
3 Iterative controller inference method
We present an approach for generating informative data sets. We first derive sufficient conditions for system inputs that lead to informative data. We then introduce an algorithm that iteratively applies the sufficient conditions to generate informative data, followed by an analysis of the convergence of the corresponding algorithm.
3.1 Iterative controller inference
By leveraging low-dimensional data informativity (Proposition 2), we generate a sequence of controllers with order $r_{1}\leq r_{2}\leq\dots\leq r_{T}$,
$\widehat{K}_{1}\in\mathbb{R}^{p\times r_{1}},\quad\widehat{K}_{2}\in\mathbb{R}^{p\times r_{2}},\quad\ldots,\quad\widehat{K}_{T}\in\mathbb{R}^{p\times r_{T}},$
which locally stabilize the system eq. 1 or eq. 2 over a sequence of nested subspaces with non-decreasing dimensions $r_{1}\leq r_{2}\leq\dots\leq r_{T}$,
$\mathcal{V}_{1}\subseteq\mathcal{V}_{2}\subseteq\dots\subseteq\mathcal{V}_{T}.$
When the dimension $r_{T}$ of $\mathcal{V}_{T}$ reaches the intrinsic dimension $n_{\min}$ of the system eq. 1 or eq. 2, then the system is stabilized in the sense of Section 2.1.
Each iteration $j=1,\dots,T$ of iterative controller inference (ICI) proceeds in two steps. In the first step of iteration $j$, the system is queried to extend the data triplet from the previous iteration $(U^{\operatorname{n}}_{j-1},X^{\operatorname{n}}_{j-1},Y^{\operatorname{n}}_{j-1})$ given by
$U^{\operatorname{n}}_{j-1}$ $=\begin{bmatrix}u^{\operatorname{n}}(t_{0})&u^{\operatorname{n}}(t_{1})&\ldots&u^{\operatorname{n}}(t_{j-1})\end{bmatrix},$
$X^{\operatorname{n}}_{j-1}$ $=\begin{bmatrix}x^{\operatorname{n}}(t_{0})&x^{\operatorname{n}}(t_{1})&\ldots&x^{\operatorname{n}}(t_{j-1})\end{bmatrix},$
$Y^{\operatorname{n}}_{j-1}$ $=\begin{bmatrix}f(x^{\operatorname{n}}(t_{0}),u^{\operatorname{n}}(t_{0}))&f(x^{\operatorname{n}}(t_{1}),u^{\operatorname{n}}(t_{1}))&\ldots&f(x^{\operatorname{n}}(t_{j-1}),u^{\operatorname{n}}(t_{j-1}))\end{bmatrix},$
to the current data triplet $(U^{\operatorname{n}}_{j},X^{\operatorname{n}}_{j},Y^{\operatorname{n}}_{j})$. To extend the data triplet, we generate an input signal $u(t_{j})$ as
$u^{\operatorname{n}}_{j-1}(t)=K_{j-1}(x^{\operatorname{n}}(t)-\bar{x})+\bar{u},$ (13)
using the controller $K_{j-1}$ from the previous iteration. Generating an input signal as in eq. 13 is motivated by a sufficient condition for system inputs to lead to informative data, which we derive in Section 3.2. Furthermore, exciting a system with an input given by eq. 13 ensures that the state trajectory will stay close to the steady state $\bar{x}$ as needed for the linearization eq. 8 of the nonlinear system eq. 1 or eq. 2 to remain meaningful. Once we have the input signal $u^{\operatorname{n}}_{j-1}(t)$ obtained via eq. 13, we query the system eq. 1 or eq. 2 to obtain its action at time $t_{j-1}$,
$y^{\operatorname{n}}(t_{j-1})=f(x^{\operatorname{n}}(t_{j-1}),u^{\operatorname{n}}(t_{j-1})).$ (14)
Then, the new data triplet $(U^{\operatorname{n}}_{j},X^{\operatorname{n}}_{j},Y^{\operatorname{n}}_{j})$ is given by the matrices
$U^{\operatorname{n}}_{j}$ $=\begin{bmatrix}U^{\operatorname{n}}_{j-1}&u^{\operatorname{n}}_{j-1}(t_{j-1})\end{bmatrix},$ (15a)
$X^{\operatorname{n}}_{j}$ $=\begin{bmatrix}X^{\operatorname{n}}_{j-1}&x^{\operatorname{n}}(t_{j-1})\end{bmatrix},$ (15b)
$Y^{\operatorname{n}}_{j}$ $=\begin{bmatrix}Y^{\operatorname{n}}_{j-1}&y^{\operatorname{n}}(t_{j-1})\end{bmatrix}.$ (15c)
Additionally, the subspace $\mathcal{V}_{j-1}$ is extended to $\mathcal{V}_{j}$ by
$\mathcal{V}_{j}=\mathcal{V}_{j-1}\cup\operatorname{span}(x^{\operatorname{n}}(t_{j-1})-\bar{x}).$
To advance the state to $x^{\operatorname{n}}(t_{j})$, we integrate the system in time using $u^{\operatorname{n}}_{j-1}(t)$ from eq. 13. Note that in the case of discrete-time systems eq. 2 and $t_{j}=t_{j-1}+1$, the system query and integration over time are equivalent such that $y^{\operatorname{n}}(t_{j-1})=x^{\operatorname{n}}(t_{j})$. In the first iteration $j=1$, the initial matrices $U^{\operatorname{n}}_{0}$, $X^{\operatorname{n}}_{0}$, and $Y^{\operatorname{n}}_{0}$ are empty, and there is no controller $K_{0}$ available. Therefore, we use a normally distributed random input signal $u(t)$ to query the system at time $t_{0}$ and to integrate the state from $x^{\operatorname{n}}(t_{0})$ to $x^{\operatorname{n}}(t_{1})$. We discuss this in more in detail in Section 3.4.
In the second step of each iteration $j$ of ICI, a new low-dimensional controller $\widehat{K}_{j}\in\mathbb{R}^{p\times r_{j}}$ corresponding to the subspace $\mathcal{V}_{j}$ is inferred with low-dimensional controller inference (see Algorithm 1) from the current data triplet $(U^{\operatorname{n}}_{j},X^{\operatorname{n}}_{j},Y^{\operatorname{n}}_{j})$. The corresponding lifted controller $K_{j}=\widehat{K}_{j}V_{j}^{\mkern-1.5mu\mathsf{T}}$ stabilizes the system eq. 1 or eq. 2 over the subspace $\mathcal{V}_{j}$, which means that it stabilizes all linearizations eq. 8 that describe the shifted and projected data
$U_{j}=U^{\operatorname{n}}_{j}-\bar{u},\quad\widehat{X}=V_{j}^{\mkern-1.5mu\mathsf{T}}(X^{\operatorname{n}}_{j}-\bar{x})\quad\text{and}\quad\widehat{Y}=V_{j}^{\mkern-1.5mu\mathsf{T}}(Y^{\operatorname{n}}_{j}-f(\bar{x},\bar{u})).$
If the dimension of $\mathcal{V}_{j}$ equals the intrinsic dimension $n_{\min}$ of the system eq. 1 or eq. 2, the algorithm terminates and returns the controller $K_{j}$, which stabilizes the system. Otherwise, the process is continued with iteration $j+1$.
3.2 Adaptive inputs for sampling informative data
The key step of the ICI method is the generation of the input signal via eq. 13. We now show under which circumstances this specific way of constructing an input signal leads to informative data. We begin by showing that informative data can be obtained only via non-zero input signals, which allows us to conclude that the input signals play an essential role in generating informative data. For the ease of exposition, we consider for now only systems with linear dynamics; however, we later apply the concept of linearization to extend these results to nonlinear systems.
Lemma 1.
Let $(U,X,Y)$ be a data triplet of a linear model eq. 8 such that $X\in\mathbb{R}^{N\times T}$ has full row rank and the system input samples are $U=0\in\mathbb{R}^{p\times T}$. If the underlying linear model is unstable, then the data triplet $(U,X,Y)$ cannot be informative for stabilization by feedback for any $T\in\mathbb{N}$.
Proof.
The number of data samples $T$ in $X$ must be at least the state-space dimension such that $T\geq N$ for $X$ to have full row rank. Then, there exists an $X^{\dagger}$, which is a right inverse of $X$ such that $XX^{\dagger}=I_{N}$. For all right inverses $X^{\dagger}$, it then holds that the product $YX^{\dagger}=J_{\operatorname{x}}XX^{\dagger}=J_{\operatorname{x}}$ has the same unstable spectrum as the system matrix $J_{\operatorname{x}}$ for any $T\geq N$. Therefore, there cannot exist a right inverse $X^{\dagger}$ such that $YX^{\dagger}$ has a stable spectrum and, following part (a) of Proposition 1, the $(U,X,Y)$ cannot be informative for stabilization for any $T\geq N$, which concludes the proof. ∎
Lemma 1 implies that there are non-zero input signals that generate informative data because otherwise stabilization and control were generally impossible. Additionally, the lemma provides a sufficient stability test for models by reversing the argument: If data $(U,X,Y)$ with a zero input signal $U$ is informative for stabilization then the original model from which the data has been sampled is stable.
It has been shown in [43, Cor. 5] that the minimal number of samples required for an informative data set is the intrinsic dimension $n_{\min}$ of the system and it has been shown that such data sets exist. The key step of the proof of [43, Cor. 5] is the construction of a suitable input signal using a stabilizing controller. We now reverse this approach, which allows us to obtain a sufficient condition on the choice of the input signal to guarantee the construction of informative data triplets. This is shown in the following lemma.
Lemma 2.
Let $f$ be the model of a linear system from which the data $(U,X,Y)$ has been sampled and let $\mathcal{V}_{\min}$ be the $n_{\min}$-dimensional subspace of all model trajectories. Also, let $X\in\mathbb{R}^{N\times T}$ have the row rank $n_{\min}$. If the input signal is chosen such that $U=KX$, where $K$ is a controller that stabilizes the system described by $f$, then the data $(U,X,Y)$ is informative for stabilization by feedback over $\mathcal{V}_{\min}$.
Proof.
Since $f$ is describes a linear model, there exist matrices $A\in\mathbb{R}^{N\times N}$ and $B\times\mathbb{R}^{N\times p}$ such that
$Y=AX+BU$
holds for the given data $(U,X,Y)$. Choosing the input samples as $U=KX$ then yields
$Y=AX+BKX=(A+BK)X.$ (16)
Because $X$ is assumed to have row rank $n_{\min}$, it follows that $T\geq n_{\min}$. Furthermore, there exists an orthogonal basis matrix $V\in\mathbb{R}^{N\times n_{\min}}$ of the subspace $\mathcal{V}_{\min}$ such that $x(t)=V\hat{x}(t)$ holds for all $t\geq 0$, where $\hat{x}(t)$ is the state of a linear model of dimension $n_{\min}$ of the form
$\dot{\hat{x}}(t)=\widehat{A}\hat{x}(t)+\widehat{B}u(t),$ (17)
with $\widehat{A}\in\mathbb{R}^{n_{\min}\times n_{\min}}$ and $\widehat{B}\in\mathbb{R}^{n_{\min}\times p}$. The system matrices in eq. 17 are obtained via projection onto $\mathcal{V}_{\min}$ using the orthogonal basis matrix $V$ such that $\widehat{A}=V^{\mkern-1.5mu\mathsf{T}}AV$ and $\widehat{B}=V^{\mkern-1.5mu\mathsf{T}}B$. Via Proposition 2, we know that $(U,X,Y)$ is informative for stabilization over $\mathcal{V}_{\min}$ if the projected data set $(U,\widehat{X},\widehat{Y})$, where $X=V\widehat{X}$ and $Y=V\widehat{Y}$, is informative for stabilization. Inserting the projected data into eq. 16 yields
$\widehat{Y}=V^{\mkern-1.5mu\mathsf{T}}V\widehat{Y}=V^{\mkern-1.5mu\mathsf{T}}Y=V^{\mkern-1.5mu\mathsf{T}}(A+BK)X=V^{\mkern-1.5mu\mathsf{T}}(A+BK)V\widehat{X}=\widehat{A}\widehat{X}+\widehat{B}\widehat{K}\widehat{X},$ (18)
where $\widehat{A}=V^{\mkern-1.5mu\mathsf{T}}AV$ and $\widehat{B}=V^{\mkern-1.5mu\mathsf{T}}B$ are as above and $\widehat{K}=KV$. Because $X$ has row rank $n_{\min}$ and $\mathcal{V}_{\min}$ contains all state trajectories of the model $f$, it follows that $\widehat{X}\in\mathbb{R}^{n_{\min}\times T}$ has full row rank $n_{\min}$ such that there exists a right inverse $\widehat{X}^{\dagger}$ of $\widehat{X}$. From eq. 18 it then follows that
$\widehat{Y}\widehat{X}^{\dagger}=\widehat{A}+\widehat{B}\widehat{K}.$ (19)
Because $\mathcal{V}_{\min}$ is the subspace of all trajectories of the model $f$, the basis matrix $V$ spans an invariant subspace of $A$ such that the spectrum of the projection is part of the spectrum of the full matrix [43]. Following the assumption that $K$ is a controller that stabilizes the system described by the model $f$, we have that the matrix
$V^{\mkern-1.5mu\mathsf{T}}(A+BK)V=\widehat{A}+\widehat{B}\widehat{K}$
is stable, too. From eq. 19, we then get that the product $\widehat{Y}\widehat{X}^{\dagger}$ is stable. By applying part (a) of Proposition 1, the projected data triplet $(U,\widehat{X},\widehat{Y})$ is informative for stabilization such that from Proposition 2, the result of this lemma follows. ∎
We now extend the result from Lemma 2 to the more general case of nonlinear systems.
Theorem 1.
Let $f$ be a model with unstable steady state $(\bar{x},\bar{u})$ from which the data $(U^{\operatorname{n}},X^{\operatorname{n}},Y^{\operatorname{n}})$ has been sampled, with the state data $X^{\operatorname{n}}=\begin{bmatrix}x^{\operatorname{n}}(t_{1})&\ldots&x^{\operatorname{n}}(t_{T})\end{bmatrix}\in\mathbb{R}^{N\times T}$ being close to the steady state such that $\lVert x^{\operatorname{n}}(t_{j})-\bar{x}\rVert\leq\epsilon$ holds for some $\epsilon>0$ small enough and $j=1,\ldots,T$. Also, let the the shifted data $\widetilde{X}$ obtained from $X^{\operatorname{n}}$ via eq. 9 have the row rank $n_{\min}$. If $U^{\operatorname{n}}$ is chosen via eq. 5 for a $K$ that stabilizes the system described by $f$ locally, then the data $(U^{\operatorname{n}},X^{\operatorname{n}},Y^{\operatorname{n}})$ is informative for local stabilization by linear feedback.
Proof.
First, we observe that for $\epsilon$ small enough, the shifted data triplet $(U,\widetilde{X},\widetilde{Y})$ obtained via eq. 9 approximates the idealized data $(U,X,Y)$ of the linearization eq. 8 arbitrarily well such that following the theory in [23, Sec. 10.1], we may assume that the data $(U,\widetilde{X},\widetilde{Y})$ can be described by the linearization eq. 8. Then, the $n_{\min}$-dimensional subspace $\mathcal{V}_{\min}$ of all trajectories of the linearization is spanned by the orthogonal basis matrix resulting from a singular value decomposition of the given data since $\widetilde{X}$ is assumed to have row rank $n_{\min}$ such that
$\begin{bmatrix}V_{\min}&V_{2}\end{bmatrix}\begin{bmatrix}\Sigma&0\\ 0&0\end{bmatrix}W=X,$
where the basis matrix $V_{\min}$ corresponds to the $n_{\min}$ nonzero singular values in $\Sigma$, and $W$ contains the right singular vectors. With the choice of $U^{\operatorname{n}}$ via eq. 5 with a locally stabilizing controller $K$, the assumptions of Lemma 2 are satisfied for the linearization eq. 8, which means that the shifted data triplet $(U,\widetilde{X},\widetilde{Y})$ is informative for stabilization over $\mathcal{V}_{\min}$. From [42, Cor. 1] it follows that if $(U,\widetilde{X},\widetilde{Y})$ is informative for stabilization, then $(U^{\operatorname{n}},X^{\operatorname{n}},Y^{\operatorname{n}})$ is informative for local stabilization over $\mathcal{V}_{\min}$, which concludes the proof. ∎
The results of Theorem 1 suggest that it is sufficient to use a stabilizing input signal for the generation of informative data. A stabilizing input signal can be obtained, for example, via feedback as $u(t)=Kx(t)$, where $K$ is a stabilizing controller. Additionally, for the use of linearizations, the sampled data has to stay close to the steady state of interest $(\bar{x},\bar{u})$ to satisfy the assumptions of Theorem 1. Therefore, dynamical instabilities that drive the system away from the steady state are undesired and need to be suppressed by a stabilizing control signal. However, Lemmas 2 and 1 assume a priori access to a stabilizing controller to generate the control signal for the trajectory samples. If such a controller would be available, there would be no need to construct a controller from data. We approach this causality dilemma via the ICI method as outlined in Section 3.1. Thereby, we use a sequence of stabilizing controllers over increasingly higher dimensional subspaces to iteratively update the input signal for the next sampling step by inferring a suitable controller.
3.3 Convergence analysis
Building on the results from Section 3.2, we now study the convergence of the ICI method as described in Section 3.1. A bound for the maximum number of iteration steps taken by ICI follows under some mild assumptions directly from the system identification theory.
Corollary 1.
Let the iteratively constructed data triplet $(U^{\operatorname{n}}_{j},X^{\operatorname{n}}_{j},Y^{\operatorname{n}}_{j})$ defined in eq. 15 from the ICI method be such that the concatenated matrix $\begin{bmatrix}\widetilde{X}_{j}^{\mkern-1.5mu\mathsf{T}}&U_{j}^{\mkern-1.5mu\mathsf{T}}\end{bmatrix}^{\mkern-1.5mu\mathsf{T}}$ has full column rank for all $j=1,\ldots,n_{\min}+p$, where $\widetilde{X}_{j}\in\mathbb{R}^{N\times j}$ and $U_{j}\in\mathbb{R}^{p\times j}$ are shifted data matrices eq. 9, and the columns of $X^{\operatorname{n}}_{j}$ are sufficiently close to the steady state $\bar{x}$ such that $\lVert x^{\operatorname{n}}(t_{j})-\bar{x}\rVert\leq\epsilon$ for $\epsilon>0$ small enough. Then, it is guaranteed that a stabilizing controller is inferred with ICI from $(U^{\operatorname{n}}_{j},X^{\operatorname{n}}_{j},Y^{\operatorname{n}}_{j})$ for $j=n_{\min}+p$.
Proof.
The result follows from the sampling number for system identification over low-dimensional subspaces [37, 43] and the rank condition on the concatenated matrix consisting of the shifted data. ∎
Corollary 1 states that under the condition that each generated data sample contains new system information in terms of the concatenated matrix rank, the ICI method terminates with a locally stabilizing controller after $n_{\min}+p$ steps. We note that $n_{\min}+p$ steps is only a worst case bound as it corresponds to the identification of the underlying linearized system. This worst case iteration bound in Corollary 1 can be improved under stricter assumptions as the following theorem shows.
Theorem 2.
Let $\mathcal{V}_{j}\subseteq\mathbb{R}^{N}$ be a $j$-dimensional subspace and let $K_{j-1}\in\mathbb{R}^{p\times N}$ be a state-feedback controller that is stabilizing the system described by the model $f$ over the subspace $\mathcal{V}_{j}$. Given the data triplet $(U^{\operatorname{n}}_{j},X^{\operatorname{n}}_{j},Y^{\operatorname{n}}_{j})$ from the ICI method as in eq. 15, we assume that the shifted state samples $\widetilde{X}_{j}\in\mathbb{R}^{N\times j}$ obtained from $X^{\operatorname{n}}_{j}$ via eq. 9 have full column rank $j$, the columns of $X^{\operatorname{n}}_{j}$ are sufficiently close to the steady state $\bar{x}$ in the sense that $\lVert x^{\operatorname{n}}(t_{j})-\bar{x}\rVert\leq\epsilon$ for $\epsilon>0$ small enough, and the input samples $U^{\operatorname{n}}_{j}$ satisfy eq. 5 columnwise with the state samples $X^{\operatorname{n}}_{j}$ and the controller $K_{j-1}$. Then, the data triplet $(U^{\operatorname{n}}_{j},X^{\operatorname{n}}_{j},Y^{\operatorname{n}}_{j})$ is informative for stabilization over $\mathcal{V}_{j}$. If $j=n_{\min}$ and $\mathcal{V}_{j}$ is so that $x^{\operatorname{n}}(t)-\bar{x}\in\mathcal{V}_{j}$ for all $x^{\operatorname{n}}(t)$ generated by the model $f$ with $\lVert x^{\operatorname{n}}(t_{j})-\bar{x}\rVert\leq\epsilon$, then the controller inferred from $(U^{\operatorname{n}}_{j},X^{\operatorname{n}}_{j},Y^{\operatorname{n}}_{j})$ is guaranteed to locally stabilize the underlying system.
Proof.
From the construction of the input samples $U^{\operatorname{n}}_{j}$ and the full column rank of the state samples $X^{\operatorname{n}}_{j}$, the conditions of Theorem 1 are satisfied for the nonlinear system projected onto the low-dimensional subspace $\mathcal{V}_{j}$. Therefore, the data triplet $(U^{\operatorname{n}}_{j},X^{\operatorname{n}}_{j},Y^{\operatorname{n}}_{j})$ must be informative for stabilization over $\mathcal{V}_{j}$. Since the controller inferred from the informative data set is stabilizing all linearizations that describe the shifted and projected data, it follows that if $j=n_{\min}$ and $\mathcal{V}_{j}$ containing all state trajectories generated by $f$, the linearization of the underlying system is stabilized by the inferred controller. Therefore, the controller is guaranteed to locally stabilize the underlying system, which concludes the proof. ∎
Let us discuss the implications of Theorem 2. First, in contrast to Theorem 1, we consider in Theorem 2 the use of low-dimensional subspaces $\mathcal{V}_{j}$, which reduce the sample complexity of the task of stabilization to the subspace dimension. Under the assumption that the subspace $\mathcal{V}_{j}$ with $j=n_{\min}$ contains all possible state trajectories of the sampled model around the steady state $\bar{x}$ and that the controller $K_{j-1}$ used for the generation of the complete input signal $U^{\operatorname{n}}_{j}$ is stabilizing over $\mathcal{V}_{j}$, Theorem 2 reduces the iterations of ICI from $n_{\min}+p$ in Corollary 1 to $n_{\min}$.
Considering the assumptions made in Theorem 2, we have to note that in the ICI method as presented in Section 3.1, the assumption that the same $K_{j-1}$ has been used for the generation of $U^{\operatorname{n}}_{j}$ will not be satisfied since the controller is updated in every step of the method. However, we provide a numerical extension of the method to satisfy this assumption with further details in Section 3.5. Also, we can see that the assumption of $K_{j-1}$ being stabilizing over $\mathcal{V}_{j}$ can be interpreted as $\mathcal{V}_{j}$ not containing more unstable dynamics than $\mathcal{V}_{j-1}$. To see this, we use from Section 3.1 the fact that in ICI, we have that $\mathcal{V}_{j-1}\subseteq\mathcal{V}_{j}$ holds for all $j\geq 1$, and we assume for simplicity of presentation that the corresponding basis matrices $V_{j-1}$ and $V_{j}$ are computed via the Gram-Schmidt procedure so that the columns of $V_{j-1}$ are the first $r_{j-1}$ many columns of $V_{j}$. For the controller obtained in the previous iteration via $K_{j-1}=\widehat{K}_{j-1}V_{j-1}^{\mkern-1.5mu\mathsf{T}}$, it follows that for all possible states from the extended subspace $x\in\mathcal{V}_{j}$ it holds
$K_{j-1}x=K_{j-1}V_{j}\hat{x}=\widehat{K}_{j-1}V_{j-1}^{\mkern-1.5mu\mathsf{T}}V_{j}\hat{x}=\begin{bmatrix}\widehat{K}_{j-1}&0\end{bmatrix}\hat{x}=\widehat{K}_{j-1}\begin{bmatrix}\hat{x}_{1}\\ \vdots\\ \hat{x}_{j-1}\end{bmatrix}+0.$
This shows that the controller $K_{j-1}$ takes the same actions for states from $\mathcal{V}_{j}$ as for states from $\mathcal{V}_{j-1}$, and if $\mathcal{V}_{j}$ does contain unstable dynamics that are not contained in $\mathcal{V}_{j-1}$, then $K_{j-1}$ cannot be stabilizing. Note that if $K_{j-1}$ is not stabilizing over $\mathcal{V}_{j}$, this does not imply that there is no $K_{j}$ stabilizing over $\mathcal{V}_{j}$ or that the corresponding data triplet $(U^{\operatorname{n}}_{j},X^{\operatorname{n}}_{j},Y^{\operatorname{n}}_{j})$ is not informative for stabilization; see the discussions in [43].
3.4 Algorithm description
1
Input: Queryable and integrable model $f$, steady state $(\bar{x},\bar{u})$, initial value $x^{\operatorname{n}}(t_{0})$, time discretization scheme $t_{0}<t_{1}<\ldots<t_{T}$.
Output: State-feedback controller $K_{T}$.
2
3Initialize $U_{0}=[\leavevmode\nobreak\ ]$, $X_{0}=[\leavevmode\nobreak\ ]$, $Y_{0}=[\leavevmode\nobreak\ ]$, $V_{0}=[\leavevmode\nobreak\ ]$ and $K_{0}=0$.
4
5for $j=1$ to $T$ do
6       if $K_{j-1}==0$ then
7             Choose a random input signal $u^{\operatorname{n}}_{j-1}(t)$ with expectation $\mathbb{E}(u^{\operatorname{n}}_{j-1}(t))=\bar{u}$.
8
9      else
10             Generate the next control signal $u^{\operatorname{n}}_{j-1}(t)=K_{j-1}(x^{\operatorname{n}}(t)-\bar{x})+\bar{u}$.
11
12       end if
13
14      Query the model $f$ at $x^{\operatorname{n}}(t_{j-1})$ and $u^{\operatorname{n}}_{j-1}(t_{j-1})$ to obtain $y^{\operatorname{n}}(t_{j-1})=f(x^{\operatorname{n}}(t_{j-1}),u^{\operatorname{n}}_{j}(t_{j-1}))$.
15
16      Integrate the model $f$ from $t_{j-1}$ to $t_{j}$ starting at $x^{\operatorname{n}}(t_{j-1})$ and using the input signal $u^{\operatorname{n}}_{j-1}(t)$ to obtain $x^{\operatorname{n}}(t_{j})$.
17
18      if $\mathcal{V}_{j-1}==\{0\}$ or $K_{j-1}$ stabilizing over $\mathcal{V}_{j-1}$ then
19             Update the projection space $\mathcal{V}_{j}$ via $V_{j}=\texttt{orth}(V_{j-1},x^{\operatorname{n}}(t_{j-1})-\bar{x})$.
20
21      else
22             Keep $V_{j}=V_{j-1}$.
23
24       end if
25
26      Update the data triplet $(U^{\operatorname{n}}_{j-1},X^{\operatorname{n}}_{j-1},Y^{\operatorname{n}}_{j-1})$ as
$U^{\operatorname{n}}_{j}$ $=\begin{bmatrix}U^{\operatorname{n}}_{j-1}&u^{\operatorname{n}}_{j-1}(t_{j-1})\end{bmatrix},$
$X^{\operatorname{n}}_{j}$ $=\begin{bmatrix}X^{\operatorname{n}}_{j-1}&x^{\operatorname{n}}(t_{j-1})\end{bmatrix},$
$Y^{\operatorname{n}}_{j}$ $=\begin{bmatrix}Y^{\operatorname{n}}_{j-1}&y^{\operatorname{n}}(t_{j-1})\end{bmatrix}.$
27
28      Infer the next controller $K_{j}$ from $(U^{\operatorname{n}}_{j},X^{\operatorname{n}}_{j},Y^{\operatorname{n}}_{j})$ and $V_{j}$ using Algorithm 1.
29
30      if no stabilizing $K_{j}$ found then
31             Overwrite data $(U^{\operatorname{n}}_{j},X^{\operatorname{n}}_{j},Y^{\operatorname{n}}_{j})$ via Algorithm 3 and update $K_{j}$.
32
33       end if
34
35 end for
Algorithm 2 Iterative controller inference (ICI).
The complete Iterative Controller Inference (ICI) method is summarized in Algorithm 2 following the concepts introduced in the previous sections.
In the first steps of the iteration, there is no stabilizing controller known yet. Thus, we propose to generate the first data samples using random inputs $u^{\operatorname{n}}(t)$. These inputs should still be centered about the steady state control $\bar{u}$ to generate meaningful data such that we assume that in expectation it holds $\mathbb{E}(u^{\operatorname{n}}(t))=\bar{u}$. Furthermore, in the first few iterations, there is no guarantee that the current data triplet $(U^{\operatorname{n}}_{j},X^{\operatorname{n}}_{j},Y^{\operatorname{n}}_{j})$ is informative for stabilization over the current subspace $\mathcal{V}_{j}$. In this case, we keep the subspace dimension constant over multiple iterations until the data triplet becomes informative. This happens eventually after $p$ additional steps under the assumption that the state samples are linearly independent since with a subspace of dimension $r_{j}$ and $r_{j}+p$ data samples, we can identify a unique linear model in $\mathcal{V}_{j}$ that describes the given data using system identification [37].
To start the ICI method, an arbitrary initial condition $x(t_{0})\in\mathcal{X}_{0}^{\operatorname{n}}$ is sufficient as long as it is close enough to the steady state $\bar{x}$ in the sense that $\lVert x(t_{0})-\bar{x}\rVert$ is small enough such that the underlying system can be described well by its linearization eq. 8. A recommended choice is the steady state itself $x^{\operatorname{n}}(t_{0})=\bar{x}$.
One motivation for the use of ICI is that trajectories constructed by the method are expected to remain in the surrounding of the steady state $\bar{x}$ in the sense that $\lVert x(t_{j})-\bar{x}\rVert$ is small for $j=0,\ldots,T$. This property is also beneficial for other numerical problems that are solved via linearizations. The following remark discusses the use of Algorithm 2 for general system identification.
Remark 1.
Data generated via a single linear stabilizing controller does not allow for linear system identification since
$\operatorname{rank}(X)=\operatorname{rank}\left(\begin{bmatrix}X\\ KX\end{bmatrix}\right)\leq N<N+p$
holds for any data matrix $X$ and fixed feedback $K$; see [37]. However, the data generated by Algorithm 2 is typically well suited nevertheless for system identification. First, only $p$ additional iteration steps without subspace extension are needed to satisfy the rank conditions of the data matrices over the low dimensional subspace. These additional steps need to generate independent data, which can typically obtained using random input signals. Second, as for the construction of controllers based on linearization theory, data must be generated close to the steady state of interest to yield accurate results in linear system identification. Therefore, the use of stabilizing controllers in the data generation results in the desired behavior of trajectories to remain in the vicinity of the steady state.
The ICI method as shown in Algorithm 2 contains a numerical extension in Algorithm 2 that aims to improve the numerical stability of the method and allows for the constructed data triplets to satisfy the assumptions made in Theorem 2. This will be discussed in the following section.
3.5 Numerical stability and re-projections
1
Input: Basis matrix $V=\begin{bmatrix}v_{1}&\ldots&v_{r}\end{bmatrix}\in\mathbb{R}^{N\times r}$, steady state $(\bar{x},\bar{u})$, queryable system $f\colon\mathbb{R}^{N}\times\mathbb{R}^{p}\to\mathbb{R}^{N}$, state scaling $\alpha>0$, number of data samples to generate $T$, optional controller $K\in\mathbb{R}^{p\times N}$.
2
Output: Re-projected data triplet $(U^{\operatorname{n}},X^{\operatorname{n}},Y^{\operatorname{n}})$.
3
4Initialize $U^{\operatorname{n}}=[\leavevmode\nobreak\ ]$, $X^{\operatorname{n}}=[\leavevmode\nobreak\ ]$, $Y^{\operatorname{n}}=[\leavevmode\nobreak\ ]$ and $k=1$.
5
6while $k\leq T$ do
7       if $k\leq r$ then
8             Normalize $x=\alpha\frac{v_{k}}{\lVert v_{k}\rVert_{2}}$.
9
10      else
11             Compute $\tilde{x}=\sum_{j=1}^{r}\beta_{j}v_{j}$, with random coefficients $\beta_{j}$.
12
13            Normalize $x=\alpha\frac{\tilde{x}}{\lVert\tilde{x}\rVert_{2}}$.
14
15       end if
16
17      if controller $K$ is given then
18             Compute the input signal $u^{\operatorname{n}}=Kx+\bar{u}$.
19
20      else
21             Choose a random input signal $u^{\operatorname{n}}$ with expectation $\mathbb{E}(u^{\operatorname{n}})=\bar{u}$.
22
23       end if
24
25      Query the system $y^{\operatorname{n}}=f(x+\bar{x},u^{\operatorname{n}})$.
26
27      Update the data matrices
$U^{\operatorname{n}}=\begin{bmatrix}U^{\operatorname{n}}&u^{\operatorname{n}}\end{bmatrix},\quad X^{\operatorname{n}}=\begin{bmatrix}X^{\operatorname{n}}&x+\bar{x}\end{bmatrix},\quad Y^{\operatorname{n}}=\begin{bmatrix}Y^{\operatorname{n}}&y^{\operatorname{n}}\end{bmatrix}.$
28
29      Increment $k\leftarrow k+1$.
30
31 end while
Algorithm 3 Data collection via re-projection with informed controls.
It has been observed in [43] that controller inference (Algorithm 1) can become numerically unstable with increasing amounts of data due to the poor conditioning of the data matrices obtained from unstable systems. The proposed remedy to the numerical instability is the use of a re-projection scheme, which improves the quality of the data used for controller inference. Algorithm 3 shows such a re-projection scheme in the context of our adaptive sampling method for nonlinear dynamical systems.
The general goal of Algorithm 3 is the construction of well-conditioned data triplets $(U^{\operatorname{n}},X^{\operatorname{n}},Y^{\operatorname{n}})$ for the use in ICI (Algorithm 2). To this end, for a given basis matrix $V\in\mathbb{R}^{N\times r}$ of an $r$-dimensional subspace $\mathcal{V}$ and a queryable function $f\colon\mathbb{R}^{N}\times\mathbb{R}^{p}\to\mathbb{R}^{N}$, the data triplet $(U^{\operatorname{n}},X^{\operatorname{n}},Y^{\operatorname{n}})$ is computed so that the first $r$ columns of $X^{\operatorname{n}}$ are linearly independent and that all columns of $X^{\operatorname{n}}-\bar{x}$ lie in $\mathcal{V}$. Additionally, if $V$ is an orthogonal basis, we have that for the projected data it holds that
$V^{\mkern-1.5mu\mathsf{T}}(X^{\operatorname{n}}-\bar{x})=\begin{bmatrix}\alpha I_{r}&\widetilde{X}_{2}\end{bmatrix}.$ (20)
The scaling parameter $\alpha>0$ is used to control the distance of the data samples $x$ from the given steady state $\bar{x}$. This is needed for the generated data to satisfy the assumption used in linearization theory, namely that $\lVert x\rVert=\lVert x^{\operatorname{n}}-\bar{x}\rVert$ is small, for $x^{\operatorname{n}}$ being the state of the nonlinear model described by $f$. When $(U^{\operatorname{n}},X^{\operatorname{n}},Y^{\operatorname{n}})$ is used in Algorithm 1, then the projected data eq. 20 appears in the linear matrix inequalities from Proposition 1 and it has been observed in [43] that the identity matrix $I_{r}$ in the first block of eq. 20 is numerically beneficial.
In the case that a controller $K$ is given, Algorithm 3 uses this controller to generate the input sequence in $U^{\operatorname{n}}$ and to sample the data for $Y^{\operatorname{n}}$. This yields further advantages and is important in the context of ICI. First, using the controller $K$ to generate the columns of $U^{\operatorname{n}}$ allows us to satisfy the conditions of Theorem 2 to provide informative data triplets. Second, if the given controller $K$ is stabilizing over the subspace $\mathcal{V}$ corresponding to the basis matrix $V$, the generated data samples in $Y^{\operatorname{n}}$ will be either close to the steady state $\bar{x}$ in the discrete-time case or will be small in norm in the continuous-time case. Therefore, the data in $Y^{\operatorname{n}}$ is suitable for linearization-based methods such as controller inference. Finally, the use of a controller $K$ in Algorithm 3 leads to the following result, which allows to potentially reduce the number of samples in the generated data triplet.
Theorem 3.
Let $f$ be a model with the steady state $(\bar{x},\bar{u})$, and let $\mathcal{V}$ be an $r$-dimensional subspace with orthogonal basis matrix $V\in\mathbb{R}^{N\times r}$ over which $K=\widehat{K}V^{\mkern-1.5mu\mathsf{T}}$ is a locally stabilizing controller. If the data triplet $(U^{\operatorname{n}},X^{\operatorname{n}},Y^{\operatorname{n}})$ is such that for the columns of $X^{\operatorname{n}}$ we have $x^{\operatorname{n}}_{j}=\alpha v_{j}+\bar{x}$ with $\alpha>0$ so that $\lVert x^{\operatorname{n}}_{j}-\bar{x}\rVert\leq\epsilon$ holds for some $\epsilon>0$ small enough and $j=1,\ldots,T$, and the columns of $U^{\operatorname{n}}$ are obtained via eq. 5, then $(U^{\operatorname{n}},X^{\operatorname{n}},Y^{\operatorname{n}})$ is an informative data set for stabilization over $\mathcal{V}$ with $T=r$ many data samples.
Proof.
By construction of the data triplet, we have that the row rank of $X^{\operatorname{n}}-\bar{x}$ is $r$. Also, for all linear systems, whose dynamics evolve in $\mathcal{V}$, we have that $n_{\min}\leq r$ holds for the intrinsic system dimension, and via the assumptions, it holds that $V^{\mkern-1.5mu\mathsf{T}}K=\widehat{K}$ is a stabilizing controller for all these systems. Therefore, we can apply Theorem 1 to the systems restricted to the subspace $\mathcal{V}$, which proofs that $(U^{\operatorname{n}},X^{\operatorname{n}},Y^{\operatorname{n}})$ is informative for local stabilization of $f$ over $\mathcal{V}$ with only $T=r$ many data samples. ∎
Theorem 3 states that if a controller, which stabilizes the underlying system over $\mathcal{V}$, has been given to the re-projection method in Algorithm 3, the re-projected data triplet will be informative for stabilization over $\mathcal{V}$ already at step $r$ of the re-projection scheme. In other words, the number of data samples in the re-projected data set can be reduced from $T\geq r$ to $r$ while preserving their informativity. This is expected to increase the numerical stability as well as to reduce the computational costs of the ICI method. However, in particular in the presence of numerical errors, the application of more than $r$ re-projection steps may be desired to further numerically stabilize the low-dimensional controller inference. Therefore, in Algorithm 3 of Algorithm 3 additional data may be generated from random samples in $\mathcal{V}$.
The re-projection method from Algorithm 3 is incorporated into ICI (Algorithm 2) as a fall back if no stabilizing controller could be inferred from the current data triplet. This is inspired by the observation that in many cases when numerical instabilities cause the controller inference via linear matrix inequalities to fail computationally, the data quality improvements via re-projection can resolve this issue. Also, note that in this setup, re-projection is not applied in every iteration step, which keeps the number of overall model queries small. To further improve efficiency, we have added a testing procedure to Algorithm 3 that employs the theory of Theorem 3. If the number of requested data samples $T$ in Algorithm 3 is larger than the current subspace dimension $r$, we apply the low-dimensional controller inference method (Algorithm 1) after the $r$-th step of Algorithm 3 to test, if there exists a stabilizing controller $K$ over $\mathcal{V}_{j+1}$. In this case, the size of the data triplet is reduced from $T$ to $r$ for the next iteration of ICI.
4 Numerical experiments
We demonstrate the proposed ICI method on a series of numerical examples. The experiments reported in the following have been run on a machine equipped with an Intel(R) Core(TM) i7-8700 CPU at 3.20 GHz and with 16 GB main memory. The algorithms are implemented in MATLAB 9.9.0.1467703 (R2020b) on CentOS Linux release 7.9.2009 (Core). For the solution of the linear matrix inequalities, the disciplined convex programming toolbox CVX version 2.2, build 1148 (62bfcca) [12, 13] is used together with MOSEK version 9.1.9 [22] as inner optimizer. The source codes, data and results of the numerical experiments are open source/open access and available at [41].
4.1 Experimental setup and summary of results
Figure 1: In all experiments, the new iterative controller inference (ICI) method needs less system queries to generate data sets informative enough for stabilization of the underlying system, with improvement factors up to $25$. In the case of the power network example, we were not able to find a setup in which classical low-dimensional controller inference with unguided data sampling was able to stabilize the system.
We compare our proposed ICI method from Algorithm 2 to the non-adaptive low-dimensional controller inference (Algorithm 1; see [43]) using either a single trajectory of data samples or multiple trajectories, where the simulations have been restarted. In either case, random input signals drawn from a normal distribution with expectation $\mathbb{E}(u^{\operatorname{n}}(t))=\bar{u}$ have been used to excite the systems for the data generation. Due to numerical instabilities using the generated raw data for controller inference, also the non-adaptive approaches have been equipped with the re-projection scheme Algorithm 3.
An overview about the different experiments for discrete-time systems is shown in Figure 1. The reported numbers of system queries include, when necessary, applications of the re-projection scheme (Algorithm 3) for any of the methods. In all experiments, the new approach needed less system queries than the state-of-the-art methods by factors ranging from $1.5$ to $25$. More details are given in the following sections describing the individual experiments. For additional experiments with continuous-time systems, the adaptive solvers from MATLAB’s ODE Suite [33] have been used for the time integration. As these do not allow for one-to-one comparisons in terms of system evaluations due to the adaptive time stepping of the integration methods, only results for the ICI method are shown in the following.
4.2 Unstable heat flow
(a) uncontrolled simulation
(b) iterative controller inference
Figure 2: Continuous-time unstable heat flow: ICI uses $28$ data samples at discrete-time points to construct a controller that quickly stabilizes the simulation of the unstable heat flow.
(a) uncontrolled simulation
(b) iterative controller inference
(c) inferred from single trajectory
(d) inferred from multiple trajectories
Figure 3: Discrete-time unstable heat flow: The new ICI approach stabilizes the system after only $14$ system evaluations, which is around $1.6$ times less evaluations needed by either of the non-adaptive approaches. Low-order controller inference from a single or four trajectories using comparable numbers of system evaluations fail to stabilize the system.
[…]
The results for the continuous-time version of the system can be seen in Figure 2. The ICI method uses $28$ samples generated via the adaptive time integration routine ode45 leading overall to around 4 02440244\,0244 024 model evaluations. The final simulations of the uncontrolled and stabilized system in Figure 2 are computed using the ode15s solver to illustrate the independence of the computed controller from the chosen time discretization scheme in the continuous-time case.
(a) small orthogonalization tolerance
(b) large orthogonalization tolerance
Figure 4: ICI convergence behavior for discrete-time unstable heat flow: The convergence of ICI is illustrated in terms of the approximation error of new states in the current projection space $\mathcal{V}_{j}$ and the relative change of the controller between steps. The peaks of the projection error indicate re-projections of the data, which are regularly needed for accurate orthogonalizations, but not if the data is more aggressively truncated with a large orthogonalization tolerance. In both cases, the distance of the trajectory to the steady state decreases, indicating stabilization.
We then used the IMEX Euler scheme with the sampling time $\tau=0.1$ to discretize the linear system with the state dynamics implicitly in time and the inputs explicitly in time. The results for the three compared methods using similar numbers of model queries are shown in Figure 3, with ICI using $14$ queries, the controller inference from a single trajectory using $14$ model queries and the controller inference from four trajectories using $16$ queries. Only ICI is able to provide a stabilizing controller for the system with this small amount of evaluations, while the other two approaches fail. By increasing the number of model evaluations, the other methods provide stabilizing controllers after $22$ queries with a single trajectory and $24$ evaluations using four trajectories.
To investigate the general convergence behavior of the proposed ICI method, we illustrate different quantities of interest appearing during the iterations including the approximation error of the next state sample in the current projection space, the relative change of the constructed controllers between the iteration steps and the distance of the generated state trajectory to the desired steady state in Figure 4. The peaks in the projection error in Figure 4(a) result from re-projections applied to the data. The results indicate that every time the projection error becomes comparably small, the data becomes numerically challenging for the matrix inequality solver leading to the re-projection method stepping in to improve the conditioning of the data. In Figure 4(b), a larger tolerance for the orthogonalization method is used such that the constructed projection space is increasing slower in terms of dimensions and the truncation of data is smaller. Therefore, the projection error has a higher minimum level and no re-projections are needed to keep the data well-conditioned and informative. Also, we observe that in both Figures 4(a) and 4(b) the distances of the generated trajectories to the steady state of interest are decreasing. If no re-projections occur as in Figure 4(b), the trajectories are monotonically decreasing. This is not the case when re-projections occur as shown in Figure 4(a). However, in this case, the controller appears to be more efficient as the distance to the steady state descreases faster.
4.3 Brazilian interconnected power system
(a) uncontrolled simulation
(b) iterative controller inference
Figure 5: Discrete-time power network: In the uncontrolled simulation, the output trajectories tend to infinity due to the instability of the power network. ICI provides a stabilizing controller after $20$ queries. The comparing low-dimensional controller inference setups did not provide a stabilizing controller within our limit of $300$ model queries.
[…]
The simulation with the controller constructed from ICI with $20$ model queries is shown in Figure 5. The low-dimensional controller inference based on a single or multiple trajectories using also $20$ model queries was not able to stabilize the system. These simulation results are not shown here but can be found in the accompanying code package [41]. In fact, we were not able to find any setup within our limit of $300$ model queries in which the non-adaptive low-dimensional controller inference was able to provide a stabilizing controller with either a single or multiple trajectories.
4.4 Tubular reactor
(a) CT: uncontrolled simulation
(b) CT: iterative controller inference
(c) DT: uncontrolled simulation
(d) DT: iterative controller inference
Figure 6: Continuous (CT) and discrete-time (DT) tubular reactor: Both in continuous as well as discrete-time, ICI constructs stabilizing controllers. In the discrete-time case, ICI performs better than the non-adaptive low-dimensional controller inference in terms of model queries by a factor of $1.5$.
(a) fine basis extension
(b) coarse basis truncation
Figure 7: ICI convergence behavior for discrete-time tubular reactor: The convergence of ICI is illustrated in terms of the approximation error of new states in the current projection space $\mathcal{V}_{j}$ and the relative change of the controller between steps. In both cases, the distance of the trajectory to the steady state decreases, indicating stabilization.
[…]
The continuous-time and discrete-time uncontrolled simulations and the results for the new ICI approach are shown in Figure 6. In the continuous-time case, for ICI we used the adaptive ode15s solver to generate data at $6$ discrete time points. The simulation of the resulting controller is shown in Figure 6(b). In the discrete-time case, ICI needed $8$ model queries to construct a stabilizing controller. The non-adaptive low-dimensional controller inference based on random inputs using the same sample number of model evaluations cannot stabilize the system. These results can be found in the accompanying code package [41]. However, increasing the number of allowed model queries to $12$ yields stabilizing controllers via a single or two trajectories used in the low-dimensional controller inference approach. This is a 1.5×1.5$\times$1.5 × increase in the number of model queries compared to ICI.
In Figure 7, we investigate the convergence behavior of the ICI method in the discrete-time case beyond the minimum amount of data needed to construct a stabilizing controller. The plots show the approximation error of the states in the constructed projection space, the relative change of the constructed controllers and the distance of the state trajectory from the desired steady state. Similar to the results shown in Section 4.2, we observe a fast convergence in terms of the projection error in the case of a small orthogonalization tolerance and the regular applications of re-projections in this case indicated by the spikes in the projection error in Figure 7(a). In the case of coarse basis extensions using a larger orthogonalization tolerance shown in Figure 7(b), there is only one re-projection needed in the beginning. Also, the distance of the state trajectory to the steady state is decreasing more consistently in Figure 7(b) than in Figure 7(a). However, the decrease is not monotonic as it was for the heat flow example (Figure 4(b)) likely due to the oscillating nature of the instabilities.
4.5 Crystal cluster growing
(a) uncontrolled simulation
(b) iterative controller inference
(c) inferred from single trajectory
(d) inferred from multiple trajectories
Figure 8: Discrete-time crystal cluster: The three clusters quickly break apart and diverge from each other in the uncontrolled simulation and only the ICI approach yields a suitable controller using $16$ model evaluations. About $\times4$ more queries than in ICI are needed using an unguided single trajectory.
[…]
The simulations of the model can be seen in Figure 8. The outputs shown are the mean velocities of the particles of the clusters and any blow-up behavior implies that the clusters are drifting apart indefinitely. ICI provides a stabilizing controller after $16$ model queries. The non-adaptive low-dimensional controller inference using the same number of queries with a single or multiple trajectories from random inputs are not stabilizing as shown in Figures 8(c) and 8(d). Using a single trajectory only, $70$ model evaluations are needed to construct a stabilizing controller, which is around $\times4$ more queries than ICI needs. Allowing data from three trajectories reduces the number of model queries for the design of a stabilizing controller to $30$, which is still around 1.8×1.8$\times$1.8 × more than needed for ICI.
4.6 Laminar flow an obstacle
(a) uncontrolled simulation
(b) iterative controller inference
Figure 9: Laminar flow behind an obstacle: The upper plots show the velocity field in the last time step and the lower plots the mean flow velocities at the end of the channel. ICI can stabilizes the system using only $10$ system queries, while for non-adaptive controller inference 25×25$\times$25 × or 12×12$\times$12 × more model queries are needed to construct stabilizing controllers, depending on the use of only a single or six state trajectories.
[…]
The uncontrolled simulation and the results for the ICI-based feedback controller are shown in Figure 9 with snapshots of the velocity fields from the last time step and trajectories of the mean velocities at the end of the channel. ICI needed only $10$ model queries to provide a stabilizing controller for the system. The non-adaptive low-dimensional controller inference destabilized the system. Instead using either $250$ model queries in a single state trajectory or $120$ for multiple trajectories allowed also with this non-adaptive approach to construct stabilizing controllers. These numbers of model queries are $25$ and $12$ times larger than for ICI, respectively.
5 Conclusions
We introduced an adaptive sampling scheme for the construction of informative data sets for the task of constructing stabilizing feedback controllers. The approach is inspired by sufficient conditions for data informativity using stabilizing control signals as system inputs. Thereby, a sequence of stabilizing controllers is constructed over nested low-dimensional subspaces based on state data samples that are obtained by exciting the system using previous stabilizing controllers. The numerical experiments that up to an order of magnitude fewer samples are necessary than with unguided data collection. The method as presented in this work solely relies on forward evaluations of the underlying model. Future work includes combining the proposed adaptive sampling approach with adjoint information, which can further reduce the required number of data samples [42]. This is of particular interest in applications, where the instabilities affect high-dimensional subspaces such as in chaotic systems.
Acknowledgments
Parts of this work were carried out while Werner was at the Courant Institute of Mathematical Sciences, New York University, USA.
Peherstorfer was partially supported by the Air Force Office of Scientific Research (AFOSR), USA, award FA9550-21-1-0222 and FA9550-24-1-0327 (Dr. Fariba Fahroo).